OpenStudy (brooklynsshaw):

The number of marbles of different colors stored in a hat is listed below: 4 red marbles 10 green marbles 7 blue marbles Without looking in the hat, Dan takes out a marble at random. He replaces the marble and then takes out another marble from the hat. What is the probability that Dan takes out a blue marble in both draws? 2 / 21 1 / 9 1 / 2 2 / 3

OpenStudy (brooklynsshaw):

I need help fast, I'm being timed!!

OpenStudy (brooklynsshaw):

@mjdennis

OpenStudy (loser66):

how many blue are there?

OpenStudy (brooklynsshaw):

Seven

OpenStudy (loser66):

how many marbles are there in total?

OpenStudy (brooklynsshaw):

21

OpenStudy (loser66):

On the first time, \(\dfrac{blue}{total}=?\)

OpenStudy (brooklynsshaw):

7/21 ?

OpenStudy (loser66):

yup =1/3 right?

OpenStudy (loser66):

Second time around, exactly 1/3 as a result Hence, 1/3 * 1/3 =?

OpenStudy (brooklynsshaw):

1/ 9 ?

OpenStudy (loser66):

yup. done

OpenStudy (brooklynsshaw):

Thank you!

OpenStudy (loser66):

np